Temperature dependence of viscosity of water is given by the following equation: ln (viscosity) = - 11.0318 + 1057.51 / (temperature + 214.624) where temperature is in OF and viscosity is in Ibm / (ft. s). (a) Consider a dilute aqueous solution of acetic acid in water. Apply EXCEL software to calculate and plot the diffusion coefficient of acetic acid in water as a function of temperature at the temperature range between 35 to 200 OF. (You need to select a constant temperature interval of 5 OF). (b) Re-solve part (a), but replace acetic acid with methanol. (c) Re-solve part (a), but replace acetic acid with acetone.
Answer #1

For this question to find the diffusion coefficient or diffusivity, we use the relation,

where DT1 and DT2 are diffusivities at T1 and T2 and and   are viscosities of water at T1 and T2.

We find the values of DT1 and at a reference temperature of 25oC or 77oF.

For acetone- water mixure, DT1 = 1.16*10-5 cm2 /s and = 1.864*10-5 lbm/ft.s

The following is a screenshot of the excel page

Similarly we can find for the other two components using initial value of diffusivity at 77OF